OpenStudy (anonymous):

Use a sum or difference formula to find the exact value of the given expression. sin π/12

OpenStudy (anonymous):

Hint: \(\dfrac{\pi}{3}-\dfrac{\pi}{4}=\dfrac{\pi}{12}\).